# From Laboratory to Real World: A New Benchmark Towards Privacy-Preserved Visible-Infrared Person Re-Identification

### v1

@@ 4.3 Qualitative Analysis
Evaluation on the CI Protocol. The results are reported on Tab. 1. It is noted that existing VI-ReID methods require paired visible and infrared images as input, making them unsuitable for addressing the modality incomplete challenge under the CI protocol. Therefore, we report our results using four classic federated learning algorithms: FedProx [19], Fednova [37], Moon [18], and FedAvg [23], each adopting different aggregation strategies. For these methods, the setting of local training is the same, conducting ResNet-50 supervised by $\mathcal{L}_{id}$ and $\mathcal{L}_{cir}$. The comparison methods after + only modify local training stage, with all other steps remaining consistent. The compared method only changes the local training stage. It is obvious that our proposed DPPT achieves consistent improvements across the board. Additionally, we modify two VI-ReID methods, i.e., AGW [44] and DNS [14], by converting its two-stream architecture into a single-stream structure, adopting our proposed sampling approach, and removing its modality information (denoted by AGW and DNS). They remain a significant performance gap compared with our DPPT.
@@
Evaluation on ES and EI Protocols. We first replicate several recent VI-ReID methods [26, 44, 49, 43, 14] to evaluate their generalization ability under the ES protocol, shown in Tab. 2. While most methods slightly outperform our baseline B (ERM with $\mathcal{L}_{id}$ and $\mathcal{L}_{cir}$), the differences are minimal, and overall rank-1 accuracy remains low. This suggests that current VI-ReID methods still have limited capability in handling unseen environments. We encourage researchers to explore the bottlenecks that limit the generalization ability of existing methods, bringing VI-ReID closer to real world and advancing the community. Moreover, we report our method under the EI protocol. Remarkably, our DPPT achieves comparable or even superior performance under EI compared to methods evaluated under ES. Moreover, The baseline results under the EI protocol (B) are similar to those under the ES protocol (B), indicating that decentralized training does not significantly impact the model’s ability to generalize to unseen environments.
@@
Table 3: Ablation studies on SYSU-MM01 and LLCM datasets. The light purple is our final choice setting.
@@
B+CA+$\mathcal{L}_{mrb}$(edu) 43.79 44.00 30.77 33.45 40.62 37.35
B+CA+$\mathcal{L}_{mrb}$(mixed) 47.82 47.55 33.94 34.41 41.56 38.09
4.4 Ablation Study
@@
Effectiveness of Each Component. To evaluate the contribution of each component, we conduct a series of ablation experiments on SYSU-MM01 [38] and LLCM [49] datasets Under CI protocol. The results are shown in Tab. 3. Baseline denotes that we employ the FedAvg [23] to train the ResNet-50 and supervised by $\mathcal{L}_{id}$ and $\mathcal{L}_{cir}$. Compared with the grayscale, CA brings a superior improvement. This is because CA preserves original information by randomly mapping one channel to the other two, whereas grayscale discards valuable color information crucial for identity discrimination. Subsequently, our proposed $\mathcal{L}_{mrb}$ further improves the performance. In addition, to explore the role of cosine similarity in Eq. 6, we substituted it with Euclidean distance (edu) and a combination of both (mixed), as detailed in the supplementary material. The ablation study under EI is shown in Tab. 2, showing consistent improvement by adding each component.
Visualization. To demonstrate the effectiveness of each component, we visualize the distribution of learned features by t-SNE [35] and intra-class and inter-class distances, as shown in Fig. 3. Compared with (a) which only uses ResNet-50 pretrained on the ImageNet-1k, the baseline improves the discriminative of identity within the modality but still exhibits a significant cross-modality gap. Subsequently, the integration of CA effectively bridges this huge discrepancy, pushing away the intra-class and inter-class distances. Despite encouraging progress, the feature distribution map shows that the similarity between different identities is relatively close, making it hard to distinguish. Finally, with the help of MRB, the modality gap is further reduced. The reason can be attributed to that MRB handles the domain shift and identity missing well, thus achieving extremely smaller intra-class and giant inter-class distances, facilitating easier distinction between identity samples.
Figure 4: The effect of hyperparameter $\lambda$ and K on SYSU-MM01.
Hyper-Parameter Ablation. The hyper-parameter ablation of $\lambda$ and Top $K$ is as shown in Fig. 4(a) and (b). The $\lambda$ is set to 1.0 and the $K$ is set to 4.
5 Conclusion
@@
This paper presents L2RW, the first privacy-preserved benchmark for VI-ReID, aims at bringing VI-ReID closer to real-world applications. To simulate privacy constraints, we introduce three protocols: camera independence (CI), entity independence (EI) and entity sharing (ES). We also propose the DPPT method, which addresses challenges under both protocols without compromising data privacy. Extensive experiments validate the feasibility of privacy-preserved VI-ReID and demonstrate the effectiveness of our method which achieves significant improvement under privacy protection conditions. Our results further show that decentralized training does not significantly affect generalization. We hope this work inspires future research in privacy-preserved VI-ReID and advances the field.
6 Acknowledgement
This work was supported by the National Natural Science Foundation of China (Grant No. U22B2062), the Research Council of Finland (former Academy of Finland) Academy Professor project EmotionAI (grants 336116, 345122, 359854), the University of Oulu & Research Council of Finland Profi 7 (grant 352788), EU HORIZON-MSCA-SE-2022 project ACMod (grant 101130271), the Finnish Doctoral Program Network in Artificial Intelligence, AI-DOC (decision number VN/3137/2024-OKM-6), and the Startup Foundation for Introducing Talent of NUIST.

### v2

@@ IV-C Quantitative Analysis for Image-based VI-ReID
Evaluation on the CI Protocol. The results are reported on Tab. II. It is noted that existing VI-ReID methods require paired visible and infrared images as input, making them unsuitable for addressing the modality incomplete challenge under the CI protocol. Therefore, we report our results using four classic federated learning algorithms: FedProx [25], Fednova [45], Moon [26], and FedAvg [24], each adopting different aggregation strategies. For these methods, the setting of local training is the same, conducting ResNet-50 supervised by $\mathcal{L}_{id}$ and $\mathcal{L}_{cir}$. The comparison methods after + only modify the local training stage, with all other steps remaining consistent. The compared method only changes the local training stage. It is obvious that our proposed I-DPPT achieves consistent improvements across the board. Additionally, we modify two VI-ReID methods, i.e., AGW [13] and DNS [16], by converting their two-stream architecture into a single-stream structure, adopting our proposed sampling approach, and removing their modality information (denoted by AGW and DNS). They remain a significant performance gap compared with our DPPT.
Furthermore, we compare our proposed I-DPPT with existing VI-ReID methods based on centralized training, as shown in Tab. III. It can be observed that a noticeable performance gap remains between decentralized and centralized training approaches. This can be attributed to two main factors. First, centralized methods typically have access to modality and camera information; Second, they often incorporate additional modules and loss functions to alleviate the cross-modality gap. However, we observe that as the amount of training data increases, the gap between decentralized training and the best centralized methods consistently narrows. To be specific, the three datasets, namely RegDB, SYSU-MM01, and LLCM, are listed in ascending order of training data size. The rank-1 accuracy gaps between our method and the best-performing centralized methods (marked underline in table) on these datasets are 34.15%, 26.53%, and 22.76%, respectively. Considering that surveillance systems in real-world scenarios typically capture massive volumes of pedestrian data, these observations are encouraging and suggest strong potential for the practical adoption of decentralized training approaches in VI-ReID.
@@
Evaluation on ES and EI Protocols. We first replicate several recent VI-ReID methods [15, 13, 3, 14, 16] to evaluate their generalization ability under the ES protocol, shown in Tab. IV. While most methods slightly outperform our baseline B (ERM with $\mathcal{L}_{id}$ and $\mathcal{L}_{cir}$), the differences are minimal, and overall rank-1 accuracy remains low. This suggests that current VI-ReID methods still have limited capability in handling unseen environments. We encourage researchers to explore the bottlenecks that limit the generalization ability of existing methods, bringing VI-ReID closer to the real world and advancing the community. Moreover, we report our method under the EI protocol. Remarkably, our I-DPPT achieves comparable or even superior performance under EI compared to methods evaluated under ES. Moreover, the baseline results under the EI protocol (B) are similar to those under the ES protocol (B), indicating that decentralized training does not significantly impact the model’s ability to generalize to unseen environments.
TABLE V: Evaluation of our V-DPPT under CI protocol. From Tab. II we can find that methods proposed after FedAvg even perform worse than FedAvg under the CI protocol. The reason is that these FL methods are general methods and ignore the unique challenges of VI-ReID. Therefore, we did not reproduce additional FL methods and adopted FedAvg as our default. The upper part of the table lists methods based on centralized training, while the lower part is our proposed methods based on decentralized training. $\triangle$ denote the retrieval accuracy gap between our V-DPPT and the existing best centralized method.
[…]
AuxNet [21] TIFS24 51.10 - - - 46.00 - 63.60 79.89 85.25 88.31 61.07 -
CAJ [14] ICCV21 56.59 73.49 79.52 84.05 41.49 - 40.49 66.79 73.32 81.16 41.46 -
CLIP-ReID [51] AAAI23 58.40 73.20 79.80 - 45.30 - 49.00 73.00 81.20 - 50.40 -
TF-CLIP [22] AAAI24 62.30 76.20 81.60 - 47.50 - 49.40 76.80 83.70 - 51.90 -
DSAF [47] TMM25 62.67 75.74 81.17 85.71 49.00 24.93 68.01 85.63 89.08 91.38 65.07 43.30
MITML [19] CVPR22 63.74 76.90 81.72 86.28 45.31 - 49.07 67.90 75.37 81.53 45.50 -
DNS [16] ECCV24 64.40 76.85 82.00 $\underline{86.30}$ 50.68 $\underline{25.94}$ $\underline{69.35}$ $\underline{86.21}$ $\underline{90.23}$ $\underline{92.34}$ 65.93 $\underline{43.98}$
SGIEL [17] CVPR23 67.65 80.30 84.73 - 52.30 - - - - - - -
VLD [23] TIFS25 $\underline{74.30}$ $\underline{85.00}$ $\underline{89.30}$ - $\underline{61.80}$ - 66.70 85.80 89.90 - $\underline{66.00}$ -
[…]
$\triangle$ - 15.97 11.88 10.29 2.36 16.81 4.84 17.24 12.84 10.73 6.13 15.30 10.27
TABLE VI: Evaluation of our V-DPPT under EI and ES protocol. The upper part of the table is under ES, while the lower part is under EI. Note that as only one dataset is available during training, we split the camera following the CI to achieve EI, and then directly test on the other dataset. H and B denote HITSZ-VCM and BUPTCampus datasets, respectively. The upper part methods adopt centralized training, while our proposed DPPT adopts decentralized training. The underlined and bold indicate the best results for both protocols, respectively. The left of $\rightarrow$ indicates seen entities, and the right is the unseen entity.
[…]
DSAF [47] TMM25 12.07 23.75 30.08 38.12 13.35 8.67 29.41 47.53 55.96 64.33 15.36 1.97
VLD [23] TIFS25 $\underline{19.35}$ $\underline{37.36}$ $\underline{46.74}$ $\underline{53.64}$ $\underline{21.15}$ $\underline{13.12}$ $\underline{40.84}$ $\underline{60.47}$ $\underline{68.19}$ $\underline{75.92}$ $\underline{28.67}$ $\underline{8.17}$
[…]
IV-D Quantitative Analysis for Video-based VI-ReID
Evaluation on the CI Protocol. As shown in Tab. II, we reproduce four classic FL algorithms [24, 25, 45, 26] for image-level VI-ReID and find that methods proposed after FedAvg even perform worse than FedAvg under the CI protocol. The reason is that these FL methods are general methods and ignore the unique challenges of VI-ReID. Therefore, we did not reproduce additional FL methods and adopted FedAvg as our default. The results are reported in Table V. Similar to prior work, existing video-based VI-ReID methods are not directly applicable under the CI protocol since their frameworks rely on centralized data access. Therefore, we compare our proposed approach under the CI setting with methods that use centralized training. Although there remains a noticeable gap from the current state-of-the-art, the performance difference is gradually narrowing, particularly when compared to image-based VI-ReID methods. Notably, our method achieves a rank-10 accuracy of 79%, indicating that the correct target appears within the top 10 retrieved sequences with a 79% probability. This substantially reduces the need for manual verification and demonstrates the practical utility of our method in real-world applications.
Moreover, it can be observed that the performance gap between decentralized and centralized training methods is smaller on video datasets compared to image datasets. Specifically, our DPPT exhibits a rank-1 gap of 15.97% with VLD on the HITSZ-VCM dataset, and 17.24% with DNS on the BUPTCampus dataset. The potential reason is that video datasets consist of sequences of pedestrian frames, where variations in background and lighting conditions are relatively limited. This consistency facilitates learning identity-preserving representations and leads to improved robustness in handling cross-modality discrepancies. Additionally, the training set of HITSZ-VCM is larger than that of BUPTCampus, and correspondingly, the gap with the best method is smaller, which aligns with the trend observed in image-based datasets. In fact, the VLD introduces text modality as auxiliary and employs the CLIP vision encoder (ViT-B-16) and text encoder, making it significantly more complex than other methods. After excluding the VLD method, our DPPT shows only a 9.32% rank-1 accuracy gap compared to the second-best method SGIEL on the HITSZ-VCM dataset. This observation further underscores the potential of decentralized training in advancing practical applications of VI-ReID.
[…]
@@ IV-D Quantitative Analysis for Video-based VI-ReID
TABLE VII: Ablation studies on SYSU-MM01 and LLCM datasets.The The last row shows our final chosen setting.
@@
TABLE VIII: Ablation studies on HITSZ-VCM and BUPTCampus datasets.The last row shows our final chosen setting.
[…]
B+CA 52.36 39.06 16.31 41.00 40.26 $27 \bigstrut$[t]
[…]
IV-E Ablation Study
@@
Effectiveness of Each Component. To evaluate the contribution of each component, we first conduct a series of ablation experiments on SYSU-MM01 [2] and LLCM [3] datasets under the CI protocol. The results are shown in Tab. VII. Baseline denotes that we employ the FedAvg [24] to train the ResNet-50 and optimized by $\mathcal{L}_{id}$ and $\mathcal{L}_{cir}$. Compared with the grayscale, Channel Augmentation (CA) [14] brings a superior improvement. This is because CA preserves original information by randomly mapping one channel to the other two, whereas grayscale discards valuable color information crucial for identity discrimination. Subsequently, our proposed $\mathcal{L}_{mrb}$, mitigates the identity missing and domain shift issues by correct the center point across clients, thus further improves the performance. The ablation study under EI is shown in Tab. IV, showing consistent improvement by adding each component. Then, we conduct ablation experiments on two public video datasets, i.e., HITSZ-VCM [19] and BUPTCampus [21] under the CI protocol. The results are reported in Tab. VIII. Unlike in image-based VI-ReID, the CA module brings little improvement in the video setting. A potential reason is that video-based VI-ReID takes sequences of pedestrian frames as input, requiring the model to first learn a temporal-invariant representation to preserve identity information before it can effectively address cross-modality discrepancies. However, our proposed MRB still yields significant performance gains, as it effectively mitigates domain shift and identity missing issues. Regardless of whether the input is image or video data, our proposed MRB consistently enhances retrieval performance, demonstrating its effectiveness and adaptability.
Figure 5: The effect of hyperparameter $\lambda$ and K on the SYSU-MM01 dataset.
Hyper-Parameter Ablation. The hyper-parameter ablation of $\lambda$ and Top $K$ is as shown in Fig. 5(a) and (b). $\lambda$ is used to control the MRB. As $\lambda$ increases, both rank-1 and mAP accuracy consistently improve, reaching the best performance when $\lambda=1.0$. The potential reason is that identity missing and domain shift are two major challenges in decentralized VI-ReID training. Our MRB is specifically designed to mitigate these issues by learn consistent representation across clients. Therefore, increasing $\lambda$ strengthens the effect of MRB, leading to significant improvements in retrieval accuracy. $K$ controls the number of features used to compute the client identity center in our MRB. It can be observed that increasing $K$ from 1 to 4 leads to consistent improvements in both rank-1 and mAP accuracy, with the best performance achieved at $K=4$. This is because the computed identity center can be viewed as a robust representation of pedestrians within each client. Through the optimization of our MRB, this representation effectively alleviates the domain shift issue. However, when $K$ continues to increase beyond 4, retrieval performance begins to decline. The client center is affected by features that are distant from the true global representation, introducing bias and hindering the alignment of optimization across different clients toward a same target. Therefore, we set $K=4$ as the optimal value.
Figure 6: The ablation study of the proposed MRB. The values in the matrix represent the Euclidean distances between the pedestrian feature centers of two clients. A smaller distance indicates less gap between clients.
Effectiveness of MRB. To evaluate the effectiveness of the proposed MRB, we perform t-SNE visualizations on the SYSU-MM01 dataset by randomly sampling 20 images of the same identity from each client. The results are shown in Fig. 6. Each entry in the matrix indicates the Euclidean distance between the pedestrian feature centers of two clients, where smaller values indicate less domain shift. In the baseline setting, the features from certain clients (e.g., green and gray) are noticeably distant from those of other clients as well as from the global feature center, indicating a severe domain shift issue. In contrast, with the introduction of MRB, the features from all clients are better aligned toward the center. The feature distributions across different clients (indicated by different colors) appear more uniform and are closer to the global center. In addition, the total Euclidean distance of all clients decreases from 3.16 to 2.62. These observations indicate that the domain shift issue is effectively mitigated. Therefore, the proposed MRB substantially mitigates the cross-modality discrepancy, resulting in notable gains in retrieval performance.
IV-F Visualization and Discussion
Figure 7: t-SNE, cosine similarity and intra-inter distances visualization of each component. The circles and triangles denote the visible and infrared modality, respectively. Different colors denote different identities. Each entry in the cosine similarity represents the average similarity between the features of one identity in the visible modality and another identity in the infrared modality. The values range from 0 to 1, where higher diagonal values indicate stronger cross-modality identity consistency.
Visualization. To demonstrate the effectiveness of each component, we visualize the distribution of learned features by t-SNE [52], cosine similarity, and intra-class and inter-class distances, as shown in Fig. 7. In Fig. 7(a), the features extracted by the ResNet-50 pretrained on ImageNet-1k exhibit a large separation between visible and infrared modalities in the t-SNE visualization. The cosine similarity matrix also shows little distinction between diagonal and off-diagonal values, and the difference between intra-class and inter-class distances is minimal. These observations indicate that the cross-modality gap is huge and the identity discriminability of model is very poor. Subsequently, with our baseline, as illustrated in Fig. 7(b), features belonging to the same identity (denoted by color) are more tightly clustered, the diagonal entries in the similarity matrix become more distinguishable, and the gap between intra-class and inter-class distances increases. This suggests that the baseline effectively reduces intra-modality variations. However, it still struggles to eliminate the cross-modality gap, resulting in limited improvement in identity discriminability. Then, by introducing the CA module (Fig. 7(c)), the modality discrepancy is further alleviated. In the t-SNE visualization, features of the same identity across modalities are more aligned, the contrast of diagonal values in the similarity matrix improves, and the intra-inter class distance gap becomes larger. Nevertheless, due to the domain shift issue, identity separability remains suboptimal. Specifically, the inter-class clusters are not well-separated in the t-SNE plot, the diagonal vs. off-diagonal contrast in the similarity matrix is still limited, and the mean intra-class distance remains relatively high. Finally, with the introduction of our proposed MRB (Fig. 7(d)), these issues are significantly mitigated. Features of the same identity are tightly grouped, and different identities become more clearly separated in the t-SNE. The diagonal values in the similarity matrix are noticeably higher than the off-diagonal entries, and the intra-class distances are further reduced. This demonstrates that MRB effectively addresses the domain shift problem, leading to enhanced identity discriminability and improved retrieval performance.
TABLE IX: Robustness experiments on the SYSU-MM01 and LLCM datasets. The Clean setting refers to the standard, unperturbed input. FGSM, PGD, BIM, MIFGSM, and TIFGSM are classic adversarial attack methods used to evaluate model vulnerability. Corruption simulates real-world conditions such as rain, snow, fog, etc. $\downarrow$ denotes the accuracy degradation caused by attacks.
[…]
Discussion on the robustness. In practical surveillance systems, adverse weather such as rain, snow, and fog is common and can degrade image quality, making cross-modality retrieval more difficult. At the same time, VI-ReID is widely used in public safety applications and is susceptible to malicious attacks, which may reduce retrieval accuracy or expose sensitive identity information. To address these concerns, we conduct experiments that simulate environmental degradation and evaluate representative adversarial attacks in order to assess robustness and quantify potential privacy risks, as shown in Tab. IX. It can be observed that our decentralized training method only experiences a slightly drop in rank-1 accuracy than the centralized training methods DNS under the BIM attack. For other attacks (i.e., FGSM, PGD, MIFGSM and TIFGSM), our methods exhibit a smaller accuracy drop compared to DNS. It is worth noting that under the clean setting, our method shows a noticeable performance gap compared to DNS. However, under the above four adversarial attacks, our method even outperforms DNS. This improvement may be attributed to the decentralized training strategy, which enables the model to learn diverse feature representations from multiple clients. Such diversity helps mitigate overfitting and enhances the model’s robustness against adversarial perturbations. Furthermore, under simulated real-world scenarios (Corruption), our method demonstrates greater robustness, with less accuracy drop than DNS. These results highlight the practical advantages of decentralized training over centralized approaches, particularly in terms of robustness under adversarial threats and challenging real-world conditions.
Figure 8: IDLG [59] gradient inversion visualization. Subject is the target pedestrian image. The recovered image is reconstructed from the subject’s gradients using iDLG. The two images to the right of the recovered image are obtained by applying a bilateral filter and a Laplacian filter, respectively. Similarity is defined as the cosine similarity between the feature vectors of the subject image and the recovered image after being processed by the encoder. It serves as an indicator of how accurately the pedestrian’s attributes have been reconstructed. The value closer to 1 indicates a higher degree of similarity between the two images.
Discussion on privacy. To evaluate the potential privacy risks of L2RW+, we utilize iDLG [59], a gradient inversion technique, to reconstruct images that simulate gradient leakage scenarios. We adopt the Adam optimizer with a learning rate of 0.01 and perform 200 iterations. The resulting image is shown as the recovered image in Fig. 8. Due to the high level of noise in the recovered image, it is difficult to visually identify meaningful content. Therefore, we further process the recovered image using a bilateral filter and a Laplacian filter, respectively, as illustrated in Fig. 8. In the results reconstructed using centralized training (DNS), clear local features consistent with the subject image can be observed. Specifically, the head region in the visible image and the leg region in the infrared image are clearly reconstructed. In contrast, under our decentralized training method, such attributes are much less distinguishable. Furthermore, we feed both the reconstructed image and the corresponding subject image into the encoder (consistent with the training setup) and compute their cosine similarity. The results show that the cosine similarities obtained under our decentralized training are consistently lower than those from DNS. This indicates that the reconstructed images are less similar to the original subject images, suggesting that our decentralized training method provides better privacy protection than centralized training. This improvement can be attributed to the nature of the training paradigm: centralized training tends to focus on identity-specific features, whereas decentralized training, influenced by multiple clients, learns more generalized representations, making sensitive attributes harder to visually recover.
Discussion on the impact of data scale. Tab. III to Tab. VI reveal a positive correlation between data scale and both the retrieval accuracy and generalization ability of decentralized training. Specifically, as shown in Tab. III and Tab. V, the rank-1 accuracy gap between DPPT and the SOTA centralized methods decreases with larger training data, in both image and video datasets. Regarding generalization, Tab. IV and Tab. VI show that under larger-scale settings such as $L+S\rightarrow R$ and $H\rightarrow B$, DPPT achieves performance comparable to centralized baselines and even performs best in the $L+S\rightarrow R$ setting. These results suggest that constructing a larger-scale VI-ReID dataset would be beneficial for advancing decentralized training in this field.
Limitations. This work revisits VI-ReID and introduces a decentralized training approach for privacy-preserved VI-ReID. However, under the CI protocol, our method still exhibits a performance gap compared to methods with fully shared data [16, 3, 47, 17, 23]. Moreover, the overall rank-1 accuracy remains relatively low under both the ES and EI protocols, indicating that generalization to unseen environments remains a fundamental challenge for VI-ReID, regardless of whether centralized or decentralized training.
Future works. In future work, we plan to extend our research in the following three directions. 1) Improving performance under different protocols. As our benchmark primarily serves an exploratory purpose, we have demonstrated the feasibility of applying decentralized training to VI-ReID. Our current method does not incorporate additional modules or strategies, which makes it highly adaptable. Moving forward, we aim to explore more advanced techniques to enhance the model’s ability to handle cross-modality discrepancies under the CI protocol and to improve its generalization to unseen domains across all protocols. 2) Constructing larger-scale VI-ReID datasets. Experimental results under all three protocols indicate that the retrieval accuracy and generalization performance of our DPPT approach improve as the training data scale increases. However, existing public datasets are still limited in size compared to the volume of data typically collected in real-world surveillance systems. Therefore, we plan to build a larger-scale VI-ReID dataset to investigate the impact of data scale on decentralized VI-ReID. 3) Removing label information. Although our current approach achieves promising results in both image- and video-based VI-ReID, it still relies on labeled data. In practice, annotating large-scale surveillance imagery is highly impractical. Thus, after improving the supervised accuracy of our method, we will explore how to perform decentralized VI-ReID in an unsupervised manner. This step is essential for making VI-ReID more applicable in real-world deployments.
V Conclusion
@@
This paper presents L2RW+, the first privacy-preserved benchmark for VI-ReID, which aims to bring VI-ReID closer to real-world applications. To simulate privacy constraints, we introduce three protocols: camera independence (CI), entity independence (EI), and entity sharing (ES). We also propose the I-DPPT and V-DPPT for image and video data, which addresses challenges under both protocols without compromising data privacy. Extensive experiments across both image- and video-level VI-ReID datasets validate the feasibility and effectiveness of decentralized training under privacy-preserving conditions. Our results further show that decentralized training exhibits better generalization than centralized training. We also reveal that scaling up the data improves both retrieval accuracy and generalization in decentralized training. These findings demonstrate that privacy-preserved VI-ReID is both effective and practical, and we hope this work inspires future research in privacy-preserved VI-ReID and advances the field.